A certain mass of air is trapped into the sealed end of a narrow tube of uniform bore by thread of mercury 15cm long. When the tube is held vertically with the closed end at the bottom, the air column is 10cm long, but when the tube is held vertically with the open end underneath, the length of air column is 12cm. Calculate (i) the atmospheric pressure (ii) the length of the air column when the tube is horizontal.
1
Expert's answer
2020-05-26T12:43:47-0400

i)

"P \n_1\n\u200b\t\n =(p+15),V \n_1\n\u200b\t\n =10A"

"P \n_2\n\u200b\t\n =(p\u221215),V \n_1\n\u200b\t\n =12A"

"(p+15)10=(p\u221215)12"

The atmospheric pressure


"p=165\text{ cm of Hg = 1650 mm of Hg}"

ii)


"(p+15)10=(p+0)l\u2192(165+15)10=165l"

"l=10.9\ cm"